The Board of XYZ Block Makers is considering a management report that outlines the following recommendations for key operations of the facility: 1. The aggregate mixing operations wishes to purchase a new auger to speed up the mixing operation. Cost $20m, annual maintenance cost 5% of asset, useful life 15 years MACRS, projected benefit undiscounted $3m per annum 2. The machinery shop has asked for a new inspection equipment. Cost $2m, annual maintenance cost 5% of asset, useful life 5 years MACRS, projected benefit undiscounted $200k per annum. 3. The finishing department indicated that moves must be made to the drying room to conform to air quality regulatory standards based on post forming treatment spraying that is required. Capex costs $15m, annual maintenance cost 5% of asset, useful life 20 years MACRS, projected benefit undiscounted $3m per annum. 4. The plant manager wants to purchase a larger more modern safe. Cost $5m, annual maintenance cost 2% of asset, useful life 20 years MACRS, projected benefit undiscounted 50k per annum. For each project there is a single course of action that is proposed. Note that the single project proposals are also independent, for there is no interrelationship among them. The general manager can decide to allocate money to some, none or all of the various project proposals. The capital budget for the company is set at 45m. the internal MARR is set at 12%. 1. Outline the problem objective, alternatives and criteria for choosing the best alternatives 2. Prepare the recommendation to the board outlining the capital budget, and cash flow projection for the array of investments 3. The board takes the decision from your recommendation to reduce the capital budget to 38m with a view that only items of priority of economic return on investment should be pursued. a. Prepare a new problem objective, alternatives and criteria for choosing the best alternatives b. Prepare the recommendation to the board outlining the capital budget, and cash flow projection for the best selection of investments justifying your recommendation.
Answer #1
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=MARR=12%=0.12
N=Year   of Cash Flow
ANNUAL NET CASH INFLOW
Project 1 -Agrregate mixing operations $2,000,000 3miliion-(5%*20million)
Project2-Machinery Shop $100,000 (200000-(5%*2million)
Project 3-Finishing Department $2,250,000 (3million-(5%*15 million)
Project4-Plant Manager -$50,000 (50000-(2%*5 million)
N Project1-Aggreegate Mixing Project 2-Machinery shop Project 3-Finishing department Project 4-Plant Manager
N CF1 PV1=CF1/((1.12^N) CF2 PV2=CF2/(1.12^N) CF3 PV3=CF3/((1.12^N) CF4 PV4=CF4/((1.12^N)
Year Cash Flow Present Value Cash Flow Present Value Cash Flow Present Value Cash Flow Present Value
0 -$20,000,000 -$20,000,000 -$2,000,000 -$2,000,000 -$15,000,000 -$15,000,000 -$5,000,000 -$5,000,000
1 $2,000,000 $1,785,714 $100,000 $89,286 $2,250,000 $2,008,929 -$50,000 -$44,643
2 $2,000,000 $1,594,388 $100,000 $79,719 $2,250,000 $1,793,686 -$50,000 -$39,860
3 $2,000,000 $1,423,560 $100,000 $71,178 $2,250,000 $1,601,506 -$50,000 -$35,589
4 $2,000,000 $1,271,036 $100,000 $63,552 $2,250,000 $1,429,916 -$50,000 -$31,776
5 $2,000,000 $1,134,854 $100,000 $56,743 $2,250,000 $1,276,710 -$50,000 -$28,371
6 $2,000,000 $1,013,262 SUM -$1,639,522 $2,250,000 $1,139,920 -$50,000 -$25,332
7 $2,000,000 $904,698 $2,250,000 $1,017,786 -$50,000 -$22,617
8 $2,000,000 $807,766 $2,250,000 $908,737 -$50,000 -$20,194
9 $2,000,000 $721,220 $2,250,000 $811,373 -$50,000 -$18,031
10 $2,000,000 $643,946 $2,250,000 $724,440 -$50,000 -$16,099
11 $2,000,000 $574,952 $2,250,000 $646,821 -$50,000 -$14,374
12 $2,000,000 $513,350 $2,250,000 $577,519 -$50,000 -$12,834
13 $2,000,000 $458,348 $2,250,000 $515,642 -$50,000 -$11,459
14 $2,000,000 $409,240 $2,250,000 $460,395 -$50,000 -$10,231
15 $2,000,000 $365,393 $2,250,000 $411,067 -$50,000 -$9,135
16 SUM -$6,378,271 $2,250,000 $367,024 -$50,000 -$8,156
17 $2,250,000 $327,700 -$50,000 -$7,282
18 $2,250,000 $292,589 -$50,000 -$6,502
19 $2,250,000 $261,240 -$50,000 -$5,805
20 $2,250,000 $233,250 -$50,000 -$5,183
NPV=Net Present Value=Sum of PVs SUM $1,806,248 SUM -$5,373,472
NPV ofProject 1 -Agrregate mixing operations -$6,378,271
NPV of Project2-Machinery Shop -$1,639,522
NPV of Project 3-Finishing Department $1,806,248
NPV ofProject4-Plant Manager -$5,373,472
Project1,2 and4 have Negative NPV
These Projects are not accepted
Only Project 3 has Positive NPV
Project 3 (Finishing Department) is accepted